# Euclid Quick Data Release (Q1)

### v1

@@ 6.7 Decomposition of the AGN Component
Given our model’s ability to classify suspected AGN, it is not unreasonable for the reader to assume that the residuals of the repainting process and the output image can independently represent the AGN and galaxy components. Although utilising inpainting for decomposition is a worthwhile future research direction, the validity of the residuals to accurately represent the AGN component for further scientific analysis have not been tested and as such cannot be used. More significant exploration of which mask to use will also be vital for such a use-case. A mask that adapts its size and shape for each input image will likely be necessary. For a deep-learning approach applying AGN decomposition on Euclid data, readers are referred to Euclid Collaboration: Margalef-Bentabol et al. (2025), hereafter called MB25, whose method provides AGN fraction () measurements that we use to analyse our AGN candidates.
The subset of Q1 used within this work and the subset tested in MB25, feature an overlap of 425 781 sources. The confusion matrix for AGN predictions between the diffusion classifier and sources with an (AGN candidate for MB25) are shown in Table 4. The majority of sources () are in agreement by both methods not to be AGN candidates. With the remaining sources, only are in agreement, with over five times more AGN candidates selected by MB25. When comparing all the AGN candidates selected using their respective subsets of data, this number reduces down to 3.5 times more candidates selected by MB25 than the diffusion model (Table 5).
The left panel of Fig. 17, shows the AGN candidates selected by the diffusion, but predicted to have low AGN fraction by the model in MB25. Whereas the right panel shows the AGN fraction of agreed sources between the two methods. Fig. 26 shows where the MB25 selections are located on the inpainting metrics. As shown in Table 4, many of the MB25 sources are not located above the diffusion threshold. Of the sources that lie above the threshold, the majority show an . This implies that the selection is more sensitive to fainter AGN.
@@
7 Summary and conclusions
In this paper, we have used a diffusion-based inpainting model to identify AGN and QSOs using VIS images from the Q1 data (Euclid Quick Release Q1 2025) from the Euclid telescope. Our approach focuses on inpainting the brightest pixels within the centre of a galaxy, and employs a novel thresholding approach based on the reconstruction errors to differentiate AGN and QSOs from the rest of the galaxy population. Using only VIS images, our method can generate a large, reliable sample of optically-selected AGN, without the requirement of AGN labels ahead of time. Utilising a standard training and inference pipeline with minimal modifications for astronomical data, our method demonstrates high recall and significant overlap with AGN selections obtained through more traditional methods, across various wavelengths including optical, near-infrared (NIR), mid-infrared, and Xray.
@@ 7.1 Considerations on training and inference costs
Our DDPM inpainting model required approximately 50 hours of training using a single Nvidia A6000 GPU, which aligns with expectations for models of similar scale and architecture. However, the inference pipeline is heavily constrained by memory rather than computation, as demonstrated by its near-linear scaling – the inpainting of 1024 and 8192 images took roughly the same amount of time when sufficient GPU memory was available. This memory constraint is due to the need to store and iteratively manipulate large intermediate tensors during inference. Whilst training, although more computationally expensive, has a more structured process where gradients are computed and accumulated in a predictable manner.
@@
With an expanded compute resource (4 Nvidia H100 GPUs), we are able to inpaint just over 2.5 million images per day. As we have shown in Sects. 4 and 2.3, our model architecture and hyperparameters remain largely unmodified from their standard configurations. Optimisations tailored explicitly for astronomical-based data, especially in regards to the noise scheduler, will significantly reduce the number of iterations and inference time required to process each image. For example, Fig. 7 shows an opportunity for a two- to four-fold speed-up in timesteps when correcting for the high number of pure Gaussian noise images alone.
@@
The results from Sect. 6.6 show that although our predictions have a significant overlap with many of the selections presented, it also highlights that the ROC classifier produces similar overlap performance. Given this boundary is formed by a simple calculation on the data, rather than requiring a complex model, it may present itself as the preferred method. However, it must be reiterated that the current pipeline has had minimum optimisations for the data and the classification task. Therefore, the ROC classifier serves as a fundamental baseline for comparison in our study. Future enhancements to the model are expected to not only improve the accuracy of identifying AGN candidates but also refine our reconstruction error maps for better analysis of prospective AGN components. The other classifiers do not offer the same potential for improvement, making the diffusion method the most promising moving forward.
Our key contributions and conclusions include the following
@@
Compared to traditional colour-selection criteria, the approach is attaining high coverage in AGN candidates while achieving near-perfect separation of galaxies not featuring an AGN component. Reliably separating these different sources allows for creation of subsets with high purity which are vital for the cosmological objectives of the Euclid mission.
@@
By demonstrating that VIS images alone can generate reliable and extensive samples of AGN and QSOs, our method showcases the impressive capabilities of Euclid’s imaging. Given the future of the mission will cover , the ability for AGN identification using only VIS allows for the possibility of some science aims to be met without the immediate need for integration of multiple data sources. This can significantly reduce time and complexity for processing and analysing the data.
@@
Although our model performs well without significant optimisation for our use case, we provide insights into the behaviour of the model given the challenges presented with astronomical imaging. We show how the assumptions of machine-learning literature can result in non-optimal training, especially when it comes to the noise scheduler and training losses. By highlighting these inefficiencies, we can see where progress can be made in both astronomy and machine-learning domains.
Because this is the first use of these pipelines for such an application, we have routinely provided guidance on improvements and next steps to refine this work. Enhancing the model’s precision, improving the efficiency of both training and inference, as well as expanding its capabilities, will only increase its impact.
In conclusion, our results demonstrate the feasibility and effectiveness of using diffusion-based inpainting for AGN and QSO identification. The ability to use VIS images exclusively for selecting these candidates, provides a potential strategic flexibility to the Euclid mission whilst aiding in reducing the reliance on complex multispectral data integration. By proving the utility of applying our pipelines to the task of AGN identification, we believe it can be expanded and can significantly contribute to a broader range of astronomical applications, such as identifying gravitational lensing or finding transient objects.
@@ Acknowledgements.
- Adam et al. (2022) Adam, A., Coogan, A., Malkin, N., et al. 2022, arXiv preprint arXiv:2211.03812
- Adam et al. (2023) Adam, A., Stone, C., Bottrell, C., et al. 2023, arXiv preprint arXiv:2311.18002
- Assef et al. (2018) Assef, R. J., Stern, D., Noirot, G., et al. 2018, ApJS, 234, 23
- Davis et al. (2018) Davis, B. L., Graham, A. W., & Cameron, E. 2018, ApJ, 869, 113
- Davis et al. (2019) Davis, B. L., Graham, A. W., & Cameron, E. 2019, ApJ, 873, 85
- DESI Collaboration et al. (2024) DESI Collaboration, Adame, A. G., Aguilar, J., et al. 2024, AJ, 168, 58
- Dhariwal & Nichol (2021) Dhariwal, P. & Nichol, A. 2021, Advances in neural information processing systems, 34, 8780
- Dressler (1989) Dressler, A. 1989, in IAU Symposium, Vol. 134, Active Galactic Nuclei, ed. D. E. Osterbrock & J. S. Miller, 217
- Euclid Collaboration: Bisigello et al. (2024) Euclid Collaboration: Bisigello, L., Massimo, M., Tortora, C., et al. 2024, A&A, 691, A1
- Euclid Collaboration: Cropper et al. (2024) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2024, A&A, accepted, arXiv:2405.13492
- Euclid Collaboration: Margalef-Bentabol et al. (2025) Euclid Collaboration: Margalef-Bentabol, B., Wang, L., La Marca, A., et al. 2025, A&A, submitted
- Euclid Collaboration: Matamoro Zatarain et al. (2025) Euclid Collaboration: Matamoro Zatarain, T., Fotopoulou, S., Ricci, F., et al. 2025, A&A, submitted
- Euclid Collaboration: Mellier et al. (2024) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2024, A&A, accepted, arXiv:2405.13491
- Euclid Collaboration: Romelli et al. (2025) Euclid Collaboration: Romelli, E., Kümmel, M., Dole, H., et al. 2025, A&A, submitted
- Euclid Collaboration: Roster et al. (2025) Euclid Collaboration: Roster, W., Salvato, M., Buchner, J., et al. 2025, A&A, submitted
- Euclid Collaboration: Walmsley et al. (2025) Euclid Collaboration: Walmsley, M., Huertas-Company, M., Quilley, L., et al. 2025, A&A, submitted
- Euclid Quick Release Q1 (2025) Euclid Quick Release Q1. 2025, https://doi.org/10.57780/esa-2853f3b
- Ferrarese & Merritt (2000) Ferrarese, L. & Merritt, D. 2000, ApJ, 539, L9
- Gaia Collaboration: Bailer-Jones et al. (2023) Gaia Collaboration: Bailer-Jones, C. A. L., Teyssier, D., Delchambre, L., et al. 2023, A&A, 674, A41
- Gaia Collaboration: Vallenari et al. (2023) Gaia Collaboration: Vallenari, A., Brown, A. G. A., Prusti, T., et al. 2023, A&A, 674, A1
- Goodfellow et al. (2014) Goodfellow, I., Pouget-Abadie, J., Mirza, M., et al. 2014, Advances in neural information processing systems, 27
- Goodfellow et al. (2020) Goodfellow, I., Pouget-Abadie, J., Mirza, M., et al. 2020, Communications of the ACM, 63, 139
- Harrison & Ramos Almeida (2024) Harrison, C. M. & Ramos Almeida, C. 2024, Galaxies, 12, 17
- Ho et al. (2020) Ho, J., Jain, A., & Abbeel, P. 2020, Advances in neural information processing systems, 33, 6840
- Karras et al. (2020) Karras, T., Laine, S., Aittala, M., et al. 2020, in IEEE/CVF conference on computer vision and pattern recognition (IEEE), 8110–8119
- Kingma (2013) Kingma, D. P. 2013, arXiv preprint arXiv:1312.6114
- Kormendy & Richstone (1995) Kormendy, J. & Richstone, D. 1995, ARA&A, 33, 581
- Kullback & Leibler (1951) Kullback, S. & Leibler, R. A. 1951, The annals of mathematical statistics, 22, 79
- Lugmayr et al. (2022) Lugmayr, A., Danelljan, M., Romero, A., et al. 2022, in IEEE/CVF conference on computer vision and pattern recognition (IEEE), 11461–11471
- Lupton et al. (2004) Lupton, R., Blanton, M. R., Fekete, G., et al. 2004, PASP, 116, 133
- Magorrian et al. (1998) Magorrian, J., Tremaine, S., Richstone, D., et al. 1998, AJ, 115, 2285
- Nichol & Dhariwal (2021) Nichol, A. Q. & Dhariwal, P. 2021, in Proceedings of Machine Learning Research, Vol. 139, Proceedings of the 38th International Conference on Machine Learning, ed. M. Meila & T. Zhang (PMLR), 8162–8171
- Pathak et al. (2016) Pathak, D., Krahenbuhl, P., Donahue, J., Darrell, T., & Efros, A. A. 2016, in IEEE conference on computer vision and pattern recognition (IEEE), 2536–2544
- Radford et al. (2021) Radford, A., Kim, J. W., Hallacy, C., et al. 2021, in Proceedings of Machine Learning Research, Vol. 139, Proceedings of the 38th International Conference on Machine Learning, ed. M. Meila & T. Zhang (PMLR), 8748–8763
- Ramesh et al. (2022) Ramesh, A., Dhariwal, P., Nichol, A., Chu, C., & Chen, M. 2022, arXiv e-prints, arXiv:2204.06125
- Remy et al. (2023) Remy, B., Lanusse, F., Jeffrey, N., et al. 2023, A&A, 672, A51
- Saharia et al. (2022) Saharia, C., Chan, W., Saxena, S., et al. 2022, in Advances in Neural Information Processing Systems, ed. S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, & A. Oh, Vol. 35 (Curran Associates, Inc.), 36479–36494
- Sahu et al. (2019) Sahu, N., Graham, A. W., & Davis, B. L. 2019, ApJ, 876, 155
- Sampson et al. (2024) Sampson, M. L., Melchior, P., Ward, C., & Birmingham, S. 2024, Astronomy and Computing, 49, 100875
- Sargsyan et al. (2023) Sargsyan, A., Navasardyan, S., Xu, X., & Shi, H. 2023, in IEEE/CVF International Conference on Computer Vision (IEEE), 7335–7345
- Smith et al. (2022) Smith, M. J., Geach, J. E., Jackson, R. A., et al. 2022, MNRAS, 511, 1808
- Spagnoletti et al. (2024) Spagnoletti, A., Boucaud, A., Huertas-Company, M., Kabalan, W., & Biswas, B. 2024, arXiv preprint arXiv:2411.19158
- Stone & Courteau (2019) Stone, C. & Courteau, S. 2019, ApJ, 882, 6
- Stone et al. (2021) Stone, C., Courteau, S., & Arora, N. 2021, ApJ, 912, 41
- Storey-Fisher et al. (2024) Storey-Fisher, K., Hogg, D. W., Rix, H.-W., et al. 2024, ApJ, 964, 69
- Thanh-Tung & Tran (2020) Thanh-Tung, H. & Tran, T. 2020, in 2020 international joint conference on neural networks (ijcnn), 1–10
- Wang et al. (2018) Wang, Y., Tao, X., Qi, X., Shen, X., & Jia, J. 2018, Advances in neural information processing systems, 31
- Yeh et al. (2017) Yeh, R. A., Chen, C., Yian Lim, T., et al. 2017, in IEEE conference on computer vision and pattern recognition (IEEE), 5485–5493
- Yu et al. (2018) Yu, J., Lin, Z., Yang, J., et al. 2018, in IEEE conference on computer vision and pattern recognition (IEEE), 5505–5514
- Zhang et al. (2023) Zhang, L., Rao, A., & Agrawala, M. 2023, in IEEE/CVF International Conference on Computer Vision (IEEE), 3836–3847
Appendix A Reconstruction metrics

### v2

@@ 5.7 Decomposition of the AGN Component
Given our model’s ability to classify suspected AGN, it is not unreasonable for the reader to assume that the residuals of the repainting process and the output image can independently represent the AGN and galaxy components. Although utilising inpainting for decomposition is a worthwhile future research direction, the validity of the residuals to accurately represent the AGN component for further scientific analysis has not been tested and, as such, cannot be used. A more significant exploration of which mask to use will also be vital for such a use case. A mask that adapts its size and shape for each input image will likely be necessary. For a deep-learning approach applying AGN decomposition on Euclid data, readers are referred to Euclid Collaboration: Margalef-Bentabol et al. (2025), hereafter called MB25, whose method provides AGN fraction () measurements that we use to analyse our AGN candidates.
The subset of the sources from the Q1 catalogue (Euclid Collaboration: Romelli et al. 2025) used within this work and the subset tested in MB25 feature an overlap of 425 781 sources. The confusion matrix for AGN predictions between the diffusion classifier and sources with an (AGN candidate for MB25) are shown in Table 4. The majority of sources () are in agreement by both methods not to be AGN candidates. With the remaining sources, only agree, with over five times more AGN candidates selected by MB25. When comparing all the AGN candidates selected using their respective subsets of data, this number reduces down to 3.5 times more candidates selected by MB25 than the diffusion model (Table 5).
The left panel of Fig. 13 shows the AGN candidates selected by the diffusion classifier but predicted to have low AGN fraction by the model in MB25. The right panel shows the AGN fraction of sources that are agreed upon between the two methods. Fig. 26 shows where the MB25 selections are located on the inpainting metrics. As shown in Table 4, many of the MB25 sources are not located above the diffusion threshold. Of the sources that lie above the threshold, the majority show an . This implies that the selection is more sensitive to fainter AGN.
@@
6 Discussion
6.1 The effect of signal-to-noise
In machine-learning research, the improvements of generative models are often demonstrated with data sets that are of high resolution and typically noise-free. This provides a level of consistency across the full data set, ensuring the benchmarks reflect the model’s performance under ideal conditions. Whether through loss functions (as discussed in Sect. 2.3) or in the practical application of the pipelines, the model’s response to one part of the data will be similar to another.
However, the difficulty in learning the dynamics of our data is not consistent throughout the images. Higher signal-to-noise ratios and lower magnitudes provide higher contrast between the source features and the background noise, enabling the model to discern these features more accurately and faster than the low S/N counterparts. In the context of diffusion models, where the aim throughout training is to learn the dynamics of the introduced noise at different timesteps, the varying S/N of the images could again cause inconsistencies. This can be seen in both the classifier results shown in Fig. 12, where the fainter objects typically have lower recall, as well as the increased variance during inpainting of fainter images shown in Fig. 10.
To explore how the noise schedule impacts the Euclid images, Figure 15 shows the variability of pixel values as they progress through timesteps. We take a row and column of pixels (a slice) through each image, bin them by S/N, and compute the median pixel value at each index within the slice. This allows us to observe how the values change across the image. Since pixels affected by the source have much higher values than those near the edges of the image, Figure 15 compares slices through the outer regions (top) with slices through the centre of the image (bottom).
In the top row, we see that low S/N images are more significantly affected by the introduction of noise, and this effect appears earlier in the scheduler. The background noise converges to the Gaussian distribution across all S/N bins between halfway () and three-quarters () of the total noise steps. In contrast, for pixels dominated by the source (bottom row), high S/N images remain relatively unaffected through much of the schedule. The brightest images do not show noticeable changes until around , and only converge to the Gaussian distribution within the final 250 steps. It is evident that both the image’s S/N level and the position of the pixels within the image determine the number of timesteps needed for the noise to resemble the target Gaussian distribution.
To visualise the true effect of this differing behaviour, Fig. 16 shows a sample of images throughout different timesteps. Each image shown is a sample from the respective S/N bins below. It is clear that the noise scheduling does not perform uniformly across the images.
To improve the performance of a diffusion model, whether to better recreate a data set or to model more complex data, the simplest approach is to increase the number of unique denoising steps. The emphasis here is on uniqueness, as simply adding more timesteps can lead to redundancy, especially in the later stages where successive steps become too similar, add little value and increase overhead. This is precisely why Nichol & Dhariwal (2021) introduced the cosine-beta scheduler (discussed in Sect. 2.4); the original linear scheduler caused images to degrade into pure noise too early into the process.
The images in Fig. 16 show that for a large fraction of our data set, the model spends the majority of its time and computational effort learning the trivial task of denoising pure Gaussian noise, which provides no useful learning signal for galaxy morphology. This is a fundamental inefficiency in the implementation of diffusion-based models that not only affects Euclid images, but will apply to any scientific data set that exhibit varying levels of S/N. Although this has implications in regards to the efficiency and optimisation of the pipeline, it highlights the bigger issue of the data quality and variety assumptions implied when using off-the-shelf deep learning methods to the complex, heterogeneous data typical in large astronomical surveys and the wider scientific community. While a detailed implementation is beyond the scope of this work, this finding highlights a critical direction for future research: the development of adaptive noise schedulers that are conditioned on image properties like S/N. Such an advance would significantly improve both the efficiency and performance of diffusion models in scientific applications.
@@ 6.2 Considerations on training and inference costs
Our DDPM inpainting model required approximately 50 hours of training using a single Nvidia A6000 GPU, which aligns with expectations for models of similar scale and architecture. However, the inference pipeline is heavily constrained by memory rather than computation, as demonstrated by its near-linear scaling – the inpainting of 1024 and 8192 images took roughly the same amount of time when sufficient GPU memory was available. This memory constraint is due to the need to store and iteratively manipulate large intermediate tensors during inference. Training, although more computationally expensive, has a more structured process where gradients are computed and accumulated in a predictable manner.
@@
With an expanded compute resource (4 Nvidia H100 GPUs), we can inpaint just over 2.5 million images per day. As we have shown in Sects. 3 and 2.4, our model architecture and hyperparameters remain largely unmodified from their standard configurations. Optimisations tailored explicitly for astronomical-based data, especially in regards to the noise scheduler discussed in Sect. 6.1, will significantly reduce the number of iterations and inference time required to process each image. For example, Fig. 16 shows an opportunity for a two- to four-fold speed-up in timesteps when correcting for the high number of pure Gaussian noise images alone.
@@
The results from Sect. 5.6 show that although our predictions have a significant overlap with many of the selections presented, it also highlights that the ROC classifier produces similar overlap performance. Given that this boundary is formed by a simple calculation from the data, rather than requiring a complex model and therefore near-zero overhead to calculate, it may present itself as the preferred method. However, the ROC metric is static and its performance is fixed by its mathematical definition. It must be reiterated that the current pipeline for the diffusion model has had minimum optimisations for the data and the classification task and so can be expected to improve with larger training sets, more advanced network architectures, or the future incorporation of multi-wavelength data as conditioning information. The ROC classifier is also limited to outputting a single number. The generative capabilities of the diffusion model, where it outputs a reconstructed image, provides the potential for a much richer set of scientific applications beyond simple classification such as decomposition as was discussed in Sect. 5.7. Therefore, although it has shown good performance, its limited approach restricts the ROC classifier to only serve as a fundamental baseline for comparison in our study.
In regards to the PSF and Srsic flux classifier, although there are no computational costs to the user, traditional parametric fitting is a serial, per-object process that becomes a bottleneck for survey-scale datasets (Tuccillo et al. 2018). Having parallelisable and scalable alternative methods, even if they feature a large one-time training cost, allow for a strategic investment as we cover more of the extragalactic sky. For AGN identification specifically, the simple parametric models may not be well suited to handle complex or merged morphologies. This could explain the worse AGN-based recall seen for this classifier in Fig. 12.
Future enhancements to the diffusion model are expected to not only improve the accuracy of identifying AGN candidates but also refine our reconstruction error maps for better analysis of prospective AGN components. The other classifiers do not offer the same potential for improvement, making the diffusion method the most promising moving forward.
7 Conclusion
In this paper, we have used a diffusion-based inpainting model to identify AGN and QSOs using VIS images from the Q1 data (Euclid Quick Release Q1 2025) from the Euclid telescope. Our approach focuses on inpainting the brightest pixels within the centre of a galaxy and employs a novel thresholding approach based on the reconstruction errors to differentiate AGN and QSOs from the rest of the galaxy population. Using only VIS images, our method can generate a large, reliable sample of optically-selected AGN without requiring AGN labels ahead of time. Utilising a standard training and inference pipeline with minimal modifications for astronomical data, our method demonstrates high recall and significant overlap with AGN candidates obtained through more traditional methods, across various wavelengths including optical, near-infrared (NIR), mid-infrared, and X-ray.
Our key contributions and conclusions include the following:
@@ 7 Conclusion
Compared to traditional colour-selection criteria, the approach attains high coverage in AGN candidates while achieving near-perfect separation of galaxies not featuring an AGN component. Reliably separating these different sources allows for the creation of subsets with high purity, which are vital for the cosmological objectives of the Euclid mission.
@@
By demonstrating that VIS images alone can generate reliable and extensive samples of AGN and QSOs, our method showcases the impressive capabilities of Euclid’s imaging. Given that the future of the mission will cover , the ability for AGN identification using only VIS allows for the possibility of some science aims to be met without the immediate need for integrating multiple data sources. This can significantly reduce the time and complexity of processing and analysing the data.
@@
Although our model performs well without significant optimisation for our use case, we provide insights into the behaviour of the model given the challenges presented with astronomical imaging. We show how the assumptions of machine-learning literature, particularly regarding fixed noise schedules, can result in non-optimal training on data with a high dynamic range and large variations in signal-to-noise. By highlighting these inefficiencies, we identify a clear direction for future progress in both astronomy and machine-learning domains through the development of adaptive, data-aware training strategies.
Because this is the first use of these pipelines for such an application, we have routinely provided guidance on improvements and next steps to refine this work. Enhancing the model’s precision, improving the efficiency of both training and inference and expanding its capabilities will only increase its impact.
In conclusion, our results demonstrate the feasibility and effectiveness of using diffusion-based inpainting for AGN and QSO identification. The ability to use VIS images exclusively for selecting these candidates provides a potential strategic flexibility to the Euclid mission while reducing reliance on complex multispectral data integration. By proving the utility of applying our pipelines to the task of AGN identification, we believe it can be expanded and can significantly contribute to a broader range of astronomical applications, such as identifying gravitational lensing or finding transient objects.
@@ Acknowledgements.
Appendix A Masks
To explore the ability of the model at recreating galaxy images, we apply a variety of different masks to a selection of unseen images. Figure 17 shows the original image on the leftmost column, followed by the regenerated images in subsequent columns. With each mask shown on the top row, the white area shows the pixels that are fixed from the original image, whereas the black areas are generated using the diffusion model.
As our model’s aim is not to generate large collections of synthetic data, the variety and scientific validity of the produced galaxies are not put under high scrutiny. Instead, we only check that there are no obvious catastrophic mistakes. In the last two columns, we can see that due to the large mask and fewer pixels to guide the reconstruction, some images produce a visible boundary around the image where the combination of fixed and generated pixels is not as seamless as the other images. Given that the masks used for AGN detection are significantly smaller than those presented in Fig. 17, the fraction of pixels that the inpainting is conditioned on will be much higher, providing tighter constraints on the output compared to these examples. Improvements to the scheduler, as discussed in Sect. 6.1, would also reduce such artifacts.
Appendix B Alternative loss
As shown in Sect. 2.3, we applied a normalisation to by dividing each image’s error by its respective maximum pixel value, which could be from a brighter external source. To visualise the impact of amending the loss function, with respect to an images brightness, Fig. 18 shows how the error from the unmodified MSE changes as the model is trained for longer. Each line represents the error for a collection of images in a random batch, with the increased transparency indicating the passage of time where the model had been trained on more epochs. The final epoch of each model is presented by a black line. The effect of removing point-like sources, as discussed in Sect. 4.1, is also explored.
The two models which feature the reduced training set (bottom row) share similar dynamics during training, regardless of the loss used. This is largely due to many point-like sources also being some of the brightest images and therefore the scale of the data set is reduced. As the most problematic data are removed, the impact of disproportionate penalisation is significantly reduced, allowing the fainter sources to be the focus during training.
Appendix C Reconstruction metrics